OpenStudy (anonymous):

Find the derivative y=x^2csc5x.

OpenStudy (mathstudent55):

\(y = \csc x\) then \(y' = -\csc x \cot x\) Use that and product rule and chain rule.

OpenStudy (mathstudent55):

and power rule.

OpenStudy (mathstudent55):

Product rule: \( y = uv\) then \(y' = uv' + vu'\)

OpenStudy (mathstudent55):

\(y=x^2\csc5x\) \(y' = x^2(\csc 5x)' + \csc 5x(x^2)'\)

OpenStudy (mathstudent55):

Can you finish it now?

OpenStudy (anonymous):

Sure. Give me a second.

OpenStudy (anonymous):

Is it x(2-5xcot(5 x))csc(5 x)?

OpenStudy (mathstudent55):

\(y' = x^2 (-5\csc 5x \cot 5x) + \csc 5x (2x)\) \(y' =-5x^2 \csc 5x \cot 5x + 2x \csc 5x\) \(y' = x \csc 5x (2-5x \cot 5x)\)

OpenStudy (mathstudent55):

Yes, that's the same I got.

OpenStudy (anonymous):

Oh ok. Thanks!

OpenStudy (mathstudent55):

You're welcome.